OpenStudy (anonymous):

list the intervals increasing, decreasing, positive and negative for this graph. y=(x^2)/(x^2-2)-1/2

OpenStudy (anonymous):

calculus?

OpenStudy (anonymous):

advanced algebra

OpenStudy (anonymous):

wow ok hmm

OpenStudy (anonymous):

first of all, does it really say \[\large y=\frac{x^2}{(x^2-2)^{-\frac{1}{2}}}\]

OpenStudy (anonymous):

yes thats what its showing on my screen lol

OpenStudy (anonymous):

we should then rewrite it as \[y=x^2\sqrt{x^2-2}\]

OpenStudy (anonymous):

okay i gotcha so we rewrite then what?

OpenStudy (anonymous):

the domain of this beast is from \((-\infty,-\sqrt2)\cup (\sqrt2,\infty)\) because you can''t take the square root of a negative number

OpenStudy (anonymous):

on the first interval it is deceasing, on the second it is increasing odd for an algebra course

OpenStudy (anonymous):

oh and it is always greater than or equal to zero

OpenStudy (anonymous):

wow thank you so much your the only person on here who has helped me with problems from now and the pass thanks!! you deserve a trophy

OpenStudy (anonymous):

yw, hope it was more or less clear